# Relativistic and Electron Correlation Effects in Static Dipole Polarizabilities for Main-Group Elements

### v1

@@ 2.5.3 Errors estimation
Based on the composite schemes, the possible error sources causing the uncertainty in $P_{\text{final}}$ in this work are similar to those proposed in Refs. 14, 35, 34:
@@
    The first one is due to the finite basis sets used for computing $P_{\text{CCSD}}$/$P_{\text{CCSD(T)}}$ and $P_{\text{MRCISD}}$, i.e., $\Delta P_{\text{basis}}$. Since the errors converge quickly, the errors in these values are taken as half of the differences of $P$ values obtained by “d-aug-dyall.v3z” and “d-aug-dyall.v4z”, corresponding to the two most diffuse exponents for the respective atomic shells in the original basis sets “dyall.v3z” and “dyall.v4z”, respectively.14 In this work, all most accurate calculations of $p$-block elements are carried out using “dyall.acv4z”, where one more diffuse but more core basis functions are added to the original “dyall.v4z” basis set. See Ref. 36 for more details. This encourages the belief that the results in this work are also closer to the complete basis sets. For simplicity, the same empirical estimate method defined in Ref. 14 is used when results with additional basis functions are available in this work. Otherwise, a conservative choice is to include the full difference between this work and the most accurate results obtained from the literature as the error.
@@
    The second error is from $\Delta P_{\text{(T)}}$, $\Delta P_{\text{T}}$, or $\Delta P_{\text{Q}}$. Half of them is also taken into account as the error as in Ref. 14.
@@
    The next one is from $\Delta_{\text{fitting}}$ due to the numerical fitting procedure. Specifically, the standard errors of the regression coefficients are evaluated using the residuals from the least-square solutions, 37 which has been implemented in Ref. 38.
@@
The uncertainty of the final $P_{\text{final}}$ values is determined by taking the quadrature of the $\Delta P_{\text{basis}}$, $\Delta P_{\text{(T)}}$, $\Delta P_{\text{Q}}$, $\Delta P_{\text{core}}$, $\Delta P_{\text{vir}}$, $\Delta P_{\text{fitting}}$, and $\Delta P_{\text{others}}$ contributions. More details on the uncertainty estimation for each element will be presented in Section 4.
@@ 3.1 Basis sets
In this work, I mainly used the Dyall quadruple-$\zeta$ family uncontracted basis sets.39, 40, 41, 42, 43, 44, 45, 46 For Group 1–2 atoms, I used ANO-RCC basis set 47, 48 instead of the Dyall basis set. The original ANO-RCC basis sets are augmented with additional functions, extending each type of function in an even-tempered manner. The exponential coefficients of the additional augmented functions were determined using the equation $\zeta_{N+1}=\zeta_{N}^{2}/\zeta_{N-1}$, where $\zeta_{N}$ and $\zeta_{N-1}$ represent the smallest exponents for each atomic shell in the default basis sets.14 These new basis sets are labeled as “s-aug-ANO-RCC” and “d-aug-ANO-RCC” for single and double augmentations, respectively. After test calculations, the “s-aug-ANO-RCC” basis set was selected for Group 1 elements. For Group 2 elements, the “s-aug-dyall.cv4z” basis set was used for Sr, Ba, and Ra, while the “d-aug-dyall.cv4z” basis set was utilized for Be, Mg, and Ca. In addition, “dyall.cv4z” and “d-aug-dyall.cv4z” are also used in test calculations to assess the error due to finite basis sets, which will be discussed later. For $p$-block elements, I used the “dyall.acv4z” basis set, which includes one additional diffuse and more core basis functions compared to the original “dyall.v4z” basis set.
@@ 3.2 Correlation level
All calculations are performed using the DIRAC18 package.49 In general, orbitals within an energy range of -20 to 25 a.u. are correlated in CC and CI calculations. The convergence criterion for the CC method was set at $10^{-10}$. In SR-CC calculations, any deviation between SRn and SRd values was observed to be less than 1% in this work. For the MRCISD calculations, the energy criterion for convergence was set to $10^{-13}$. Further computational details for $p$-block open-shell elements are provided in the following subsections.
@@ 3.2.1 Group 13 elements
DHF orbitals, used for DC-CC calculations, are optimized by averaging over various $np$ occupations with one open-shell electron. The open shells are restricted to the $np_{1/2}$ spinors, leading to a single determinant reference as detailed in previous work 8. Additionally, DC-CI calculations were employed to obtain excited-state components, with DHF orbitals constructed from averaging over all $np_{1/2}$ and $np_{3/2}$ spinors.
@@ 4 Results
4.1 Group 1 elements
4.1.1 Comparison with literature
Table 1 summarizes the CCSD(T) results for Group 1 elements, compared with the recommended values from Ref.3. Additional results can be found in Table S1. The results obtained by Eq. (15) can be found in Table S9, and the corresponding summary of the most accurate results is listed in Table S17.
Table 1: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 1 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
Li Na K Rb Cs Fr
$\hat{H}$ State Method
@@ 4.1.1 Comparison with literature
Ref. 3 --- - --- - $164.11$ $162.7\pm 0.5$ $289.7\pm 0.3$ $319.8\pm 0.3$ $400.9\pm 0.7$ $317.8\pm 2.4$
In addition, the most accurate theoretical and experimental values for the entire periodic table have been summarized in Refs. 3 and 13. These values for Group 1 elements are collected in Table S25 to validate the results in this work. Several observations can be made and discussed as follows. In this work, only computational and experimental values close to the recommended values in Ref. 3 are compared.
The DC result ($164.31\pm 0.01$) for lithium (Li, $Z=3$) aligns closely with experimental results ($164.0\pm 3.4$ 50 and $164.2\pm 1.1$ 51). The most accurate computational value ($164.1125\pm 0.0005$ 52) is taken as the recommended value in Ref. 3. The difference between the recommended value and the result in this work is only 0.12%, which could be attributed to the absence of other relativistic effects, such as Breit or QED. It should be noted that the full relativistic contribution to $\alpha$ is $-0.07$ a.u. at both the DHF and CCSD levels, as shown in Table S1, which agrees with the value ($-0.07$ a.u.) obtained by Lim et al. at the DHF level.53
For sodium (Na, $Z=11$), the deviation between the DC result ($163.27\pm 0.09$) and the recommended value ($162.7\pm 0.5$) in Ref.3, based on the experimental result by Ekstrom et al., is less than 0.36%.54 Notably, the lower bound of the DC result is slightly larger than the upper bound of both the recommended value and another experimental value ($162.7\pm 0.1$) reported by Holmgren et al.55 However, the DC result for Na ($163.27\pm 0.09$) is close to the experimental value ($161\pm 7.5$) reported by Ma et al.56 Additionally, the DC result is slightly larger than the recommended computational value (162.88) in Refs. 57, 58, obtained using a composite scheme by Thakkar and Lupinetti. The NR and SR values of Na (163.98 and 162.99, respectively) in this work are in good agreement with the computational values of 163.90 and 162.90, respectively, as reported by Thakkar and Lupinetti.57, 58 Therefore, the primary difference between the DC value and the final recommended value in Refs. 57, 58 lies in the method of incorporating the SOC contribution to $\alpha$. Specifically, in Refs. 57, 58, a value of $-0.02$ a.u. evaluated at the DHF level from Ref. 53 is used as the SOC contribution and is added to the SR value. In contrast, this work evaluates the SOC contribution as $+0.28$ a.u. at the CCSD(T) level. It should be noted that in this study, the SOC contributions at the DHF and CCSD levels are $-0.04$ and $-0.03$ a.u., respectively. Furthermore, this study shows that the NR value also agrees well with the recent non-relativistic CCSD(T) value obtained by Smiałkowski and Tomza (163.9 a.u.) at CCSD(T)/aug-cc-pwCV5Z,59 suggesting the reliability of the method used in this work.
For K, the lower bound of the DC value ($290.64\pm 0.24$) is slightly higher than the upper bound of the recommended value ($289.7\pm 0.3$),3 which is based on experimental values.60, 61 However, the DC CCSD(T) value obtained using Eq. (15) is $289.40\pm 0.20$, as shown in Table S17, which agrees well with the recommended value in Ref. 3. Therefore, one reason for the discrepancy could be that a negative $\gamma$ obtained from Eq. (14) is invalid, which is observed when $\tau>0.1\%$ is used in Eq. (16).
Moreover, the DC value ($290.64\pm 0.24$) is in good agreement with other experimental values ($292.9\pm 6.1$ 50 and $290.6\pm 1.4$ 55). Compared to the computational values, the DC result is slightly higher than the RLCCSD value (289.1),62 where the linear approximation CCSD method is used. Additionally, the DC value ($290.64\pm 0.24$) agrees well with the recent result ($290.30\pm 0.23$) computed by the relativistic all-order method.63
The SR result (289.96) is in good agreement with the DK-CCSD(T) result ($291.1\pm 1.5$) reported by Lim et al. using all-electron calculations.64 Furthermore, the SR value (289.96) is also close to the recent result (289.6) computed by CCSD(T) with relativistic energy-consistent pseudopotentials (ECP).59
The most accurate DC value for Rb in this work is $320.08\pm 0.29$ a.u., which is in good agreement with previous experiments ($319\pm 6$ 50, $318.8\pm 1.4$ 55, $319.8\pm 0.3$ 60, 61, $319.2\pm 6.1$ 57) and the recommended value ($319.8\pm 0.3$).3 The DC value is generally higher than recent values (318.5±0.6plus-or-minus318.50.6318.5$\pm$ 0.6318.5 ± 0.665, 318.38±0.38plus-or-minus318.380.38318.38$\pm$ 0.38318.38 ± 0.3863) obtained by the relativistic all-order method and its variants, but it is in good agreement with the latest result ($319.5\pm 1.5$)66, which includes other relativistic effects, e.g., Breit and QED, etc. It should be noted that the DC value obtained using Eq. (15) is $319.08\pm 0.17$, which is less than the recommended value in Ref. 3. This suggests that a more accurate $\gamma$ should be used to obtain $\alpha$ for Rb. Moreover, the SR value ($319.34\pm 0.01$) is slightly larger than the latest result (317.4) computed by scalar-relativistic CCSD(T) with the ECP method.59
The DC value for Cs ($398.24\pm 0.34$) is slightly lower than the value from previous theoretical work, which was 399.0, reported by Borschevsky et al.,4 computed using the CCSD(T) method with the Fægri basis set.67 The discrepancy could be attributed to the different basis set and fitting procedure used in this study. To study the impact of basis sets, I used the same Fægri set as in Ref. 4. The virtual orbitals with energies above 40 a.u. were omitted instead of 50 a.u. in the previous work to reduce computational cost. It should be noted that five data points were used to fit $\alpha$ defined in Eq. (14) using the least-squares procedure, instead of the three data points used with numerical differentiation in Ref. 4. The result ($399.52\pm 0.62$ a.u.), as shown in Table S1, was closer to 399.0 in Ref. 4. It is also in excellent agreement with the recent results (399.9±0.6plus-or-minus399.90.6399.9$\pm$ 0.6399.9 ± 0.665 and 399.74±0.55plus-or-minus399.740.55399.74$\pm$ 0.55399.74 ± 0.5563) computed by relativistic all-order methods. When Eq. (15) is used, the results from the least-squares procedure are $400.51\pm 0.2$ and $400.60\pm 0.37$ a.u. for the Fægri and s-aug-ANO-RCC basis sets, respectively. These results are closer to the recommended value ($400.9\pm 0.7$ a.u.) from Ref. 3 compared to those obtained by fitting Eq. (14). The larger difference in DC CCSD(T) values obtained with s-aug-ANO-RCC between using Eqs. (14) and (15) can be attributed to a larger $\gamma$ obtained compared to the Fægri basis. Therefore, simply removing $\gamma$ and higher-order contributions in Eq. (15) could lead to a higher numerical error for Cs. Surprisingly, this error is even as significant as the SOC contribution to $\alpha$, as shown in Table 1. The SR value ($397.40\pm 0.02$) in this work is much higher than the result (391) obtained by scalar-relativistic CCSD(T) with the ECP method.59 One reason could be the pseudopotential used to account for the scalar-relativistic effect in Ref. 59.
The DC value of Fr ($317.40\pm 0.57$) is in excellent agreement with the recommended value ($317.8\pm 2.4$) reported in Ref. 3 and the recent computational values ($317.1\pm 1.3$,7 $316.6\pm 2.4$,68 316.6±1.5plus-or-minus316.61.5316.6$\pm$ 1.5316.6 ± 1.563) using the relativistic all-order method and its variants. There is a significant difference between the results of the current work and those reported in Ref. 4, where the Fægri basis set and fewer data points were used. To investigate this discrepancy, I chose the same data points as in Ref. 4 and fitted Eq. 15. The resulting value is 318.34, which is still much larger than the value reported in Ref. 4. Therefore, the primary reason for this discrepancy could be the different basis functions used in the two studies.
4.1.2 Uncertainty estimation
Next, the uncertainty excluding $\Delta P_{\text{fitting}}$ is estimated for all Group 1 elements. In general, the errors due to $\Delta P_{\text{core}}$ and $\Delta P_{\text{vir}}$ are not estimated explicitly. They are included implicitly in the error due to $\Delta P_{\text{basis}}$. Several SR test calculations were performed for basis sets using ANO-RCC and d-aug-ANO-RCC, compared to the results with s-aug-ANO-RCC.
In SR-CC calculations of Li, as shown in Table S1, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 164.30, 164.29, and 164.28 a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.01 a.u., with half of it taken as the error due to $\Delta P_{\text{basis}}$. With the s-aug-ANO-RCC basis set, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in DC-CC calculations is 0.07 a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$. To be conservative, I take the difference between this study and the recommended value as the error, i.e., 0.20 a.u., which is mainly due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 0.21 a.u. in this study for Li is acceptable.
In SR-CC calculations of Na, as shown in Table S1, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 163.00, 162.99, and 162.97 a.u., respectively, suggesting that the error due to using s-aug-ANO-RCC basis set is about 0.02 a.u., with half of it taken as the error due to $\Delta P_{\text{basis}}$. With the s-aug-ANO-RCC basis set, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in DC-CC calculations is 1.11 a.u., with half of it taken as the error due to $\Delta P_{\text{(T)}}$. To be conservative, the difference between this study and the recommended value is taken as the error, i.e., 0.57 a.u., which is mainly due to neglecting other relativistic and higher-order correlation effects, i.e., $\Delta P_{\text{others}}$. In conclusion, the error of 0.80 a.u. in this study for Na is acceptable.
In SR-CC calculations of K, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 284.74, 289.96, and 290.93 a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.97. To be conservative, it is taken as the error due to $\Delta P_{\text{basis}}$. With the s-aug-ANO-RCC basis set, the difference ($\Delta P_{\text{(T)}}$) between CCSD and CCSD(T) values in DC-CC calculations is 12.19 a.u., which is much higher than the difference between the recommended value and the DC result in this work, i.e., 0.94 a.u. To be conservative, I take twice the latter difference as the error due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 2.12 a.u. in this study for K is acceptable.
In SR-CC calculations of Rb, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are 314.32, 319.34, and 319.89 a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.55 a.u., which is therefore taken as the error due to finite basis sets. To be conservative, I take twice the difference between the recommended value and the DC results in this work, 0.56 a.u. in total, as the error due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 0.78 a.u. in this study for Rb is acceptable.
For Cs, I performed DC-CC calculations with d-aug-ANO-RCC, and the CCSD(T) result is $398.16\pm 0.33$, which agrees well with the value evaluated at the CCSD(T)/s-aug-ANO-RCC level ($398.24\pm 0.34$). The difference of 0.08 between these two basis sets suggests that the result is converged with the ANO-RCC basis set for $\alpha$ calculations. I therefore take the difference between the results with s-aug-ANO-RCC and Fægri basis sets as the error due to the finite basis sets ($\Delta P_{\text{basis}}$), i.e., 1.28 a.u. Based on the calculations with different fitting procedures used, the numerical errors of 2.35 a.u. are added as the error due to other effects. In conclusion, the error of 2.68 a.u. in this study for Cs is acceptable.
In SR-CC calculations of Fr, the results for CCSD(T) with ANO-RCC, s-aug-ANO-RCC, and d-aug-ANO-RCC are $318.72$, $319.36$, and $319.21$ a.u., respectively, suggesting that the error due to using the s-aug-ANO-RCC basis set is about 0.15 a.u., which is taken as the error due to finite basis sets. To be conservative, I take twice the difference between the recommended value and the DC results in this work (0.40 a.u.) as the error due to neglecting other relativistic and higher-order correlation effects ($\Delta P_{\text{others}}$). In conclusion, the error of 0.81 a.u. in this study for Fr is acceptable. The recommended values, obtained from the most accurate calculations with total uncertainty, including the corresponding $\Delta P_{\text{fitting}}$, are listed in Table 1. It should be noted that for K, the DC CCSD(T) result obtained using Eq. (15) is taken as the recommended value, as shown in Table S17.
4.1.3 The correlation and relativistic effects on polarizabilities
Figure 1(a) demonstrates that non-relativistic polarizabilities increase monotonically with atomic numbers, starting from Na. In contrast, the relativistic data increase with increasing atomic numbers from Na to Cs, but decrease from Cs to Fr, with the value for Fr being close to that of Rb. This trend confirms the irregular progression of values with atomic numbers for Group 1 elements, driven by the relativistic contraction of the valence $s$ shell.53, 64 The polarizability of Na is slightly lower than that of Li due to reduced screening of the nuclear charge by the $p$ electrons in Na, resulting in more compact valence shells and lower polarizabilities.
Figure 1(b) indicates that the scalar-relativistic effect is more pronounced than the SOC effect, leading to a roughly $Z^{2}$ increase in relativistic effects on dipole polarizabilities due to the scalar-relativistic contribution. The SOC contribution to dipole polarizabilities for Group 1 atoms is small; generally, the SOC effect slightly increases the polarizabilities of Group 1 elements, except for Fr, where the SOC effect reduces dipole polarizability by nearly 2. This observation aligns with previous findings.53, 64
4.2 Group 2 elements
4.2.1 Comparison with literature
Table 2 summarizes the CCSD(T) results for Group 2 elements, compared with the recommended values from Ref. 3. Additional results can be found in Table S2. The results obtained by Eq. (15) can be found in Table S10, and the corresponding summary of the most accurate results is listed in Table S18.
Table 2: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for Group 2 elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 3.
Be Mg Ca Sr Ba Ra
$\hat{H}$ State Method

### v2

@@ 2.5.3 Errors estimation
Based on the composite schemes, the possible error sources causing the uncertainty in $P_{\text{final}}$ in this work are similar to those proposed in Refs. 20, 4, 40:
@@
    The first one is due to the finite basis sets used for computing $P_{\text{CCSD}}$/$P_{\text{CCSD(T)}}$ and $P_{\text{MRCISD}}$, i.e., $\Delta P_{\text{basis}}$. Since the errors converge quickly, the errors in these values are taken as half of the differences of $P$ values obtained by “d-aug-dyall.v3z” and “d-aug-dyall.v4z”, corresponding to the two most diffuse exponents for the respective atomic shells in the original basis sets “dyall.v3z” and “dyall.v4z”, respectively.20 In this work, all most accurate calculations of $p$-block elements are carried out using “dyall.acv4z”, where one more diffuse but more core basis functions are added to the original “dyall.v4z” basis set. See Ref. 41 for more details. This encourages the belief that the results in this work are also closer to the complete basis sets. For simplicity, the same empirical estimate method defined in Ref. 20 is used when results with additional basis functions are available in this work. Otherwise, a conservative choice is to include the full difference between this work and the most accurate results obtained from the literature as the error.
@@
    The second error is from $\Delta P_{\text{(T)}}$, $\Delta P_{\text{T}}$, or $\Delta P_{\text{Q}}$. Half of them is also taken into account as the error as in Ref. 20.
@@
    The next one is from $\Delta_{\text{fitting}}$ due to the numerical fitting procedure. Specifically, the standard errors of the regression coefficients are evaluated using the residuals from the least-square solutions, 42 which has been implemented in Ref. 43.
@@
The uncertainty of the final $P_{\text{final}}$ values is determined by taking the quadrature of the $\Delta P_{\text{basis}}$, $\Delta P_{\text{(T)}}$, $\Delta P_{\text{Q}}$, $\Delta P_{\text{core}}$, $\Delta P_{\text{vir}}$, $\Delta P_{\text{fitting}}$, and $\Delta P_{\text{others}}$ contributions.
@@ 3.1 Basis sets
In this work, I mainly used the Dyall quadruple-$\zeta$ family uncontracted basis sets.44, 45, 46, 47, 48, 49, 50, 51 For Group 1–2 atoms, I used ANO-RCC basis set 52, 53 instead of the Dyall basis set. The original ANO-RCC basis sets are augmented with additional functions, extending each type of function in an even-tempered manner. The exponential coefficients of the additional augmented functions were determined using the equation $\zeta_{N+1}=\zeta_{N}^{2}/\zeta_{N-1}$, where $\zeta_{N}$ and $\zeta_{N-1}$ represent the smallest exponents for each atomic shell in the default basis sets.20 These new basis sets are labeled as “s-aug-ANO-RCC” and “d-aug-ANO-RCC” for single and double augmentations, respectively. After test calculations, the “s-aug-ANO-RCC” basis set was selected for Group 1 elements. For Group 2 elements, the “s-aug-dyall.cv4z” basis set was used for Sr, Ba, and Ra, while the “d-aug-dyall.cv4z” basis set was utilized for Be, Mg, and Ca. In addition, “dyall.cv4z” and “d-aug-dyall.cv4z” are also used in test calculations to assess the error due to finite basis sets, which will be discussed later. For $p$-block elements, I used the “dyall.acv4z” basis set, which includes one additional diffuse and more core basis functions compared to the original “dyall.v4z” basis set.
@@ 3.2 Correlation level
All calculations are performed using the DIRAC18 package.54 In general, orbitals within an energy range of -20 to 25 a.u. are correlated in CC and CI calculations. The convergence criterion for the CC method was set at $10^{-10}$. In SR-CC calculations, any deviation between SRn and SRd values was observed to be less than 1% in this work. For the MRCISD calculations, the energy criterion for convergence was set to $10^{-13}$. Further computational details for $p$-block open-shell elements are provided in the following subsections.
@@ 3.2.1 Group 13 elements
DHF orbitals, used for DC-CC calculations, are optimized by averaging over various $np$ occupations with one open-shell electron. The open shells are restricted to the $np_{1/2}$ spinors, leading to a single determinant reference as detailed in previous work 14. Additionally, DC-CI calculations were employed to obtain excited-state components, with DHF orbitals constructed from averaging over all $np_{1/2}$ and $np_{3/2}$ spinors.
@@ 4 Results
LABEL:tab:dipole_summary summarizes the CCSD(T) and MRCISD results for main-group elements, considering non-relativistic, scalar-relativistic, and full Dirac-Coulomb relativistic effects, and compares them with the recommended values from Ref. 1. Additionally, component-resolved results are presented for the ${{}^{2}}P_{3/2}$ state of Group 13 elements, the ${{}^{4}}S_{3/2}$ state of Group 15 elements, the ${{}^{3}}P_{2}$ state of Group 16 elements, and the ${{}^{2}}P_{1/2}$ state of Group 17 elements. The corresponding results derived from Eq. (15) are available in Table S17 in the Supporting Information.
Next, only a summary of the comparison with the recommended values from Ref. 1 is presented. A detailed comparison with other computational and experimental values from the literature can also be found in the Supporting Information. Moreover, the final uncertainty for each atom is provided in LABEL:tab:dipole_summary, with a detailed uncertainty estimation for each atom available in the Supporting Information.
Table 1: Static dipole polarizabilities (in a.u.) with non-relativistic (NR), scalar-relativistic (SR), and full Dirac-Coulomb (DC) relativistic effects for main-group elements. The error due to the numerical fitting procedure ($\Delta P_{\text{fitting}}$) is included as the error bar. The recommended values (Rec.), including the uncertainty estimation as the error bar, are also listed and compared to the counterparts from Ref. 1.
$\hat{H}$ State Method Period
1 2 3 4 5 6 7
Group 1 Li Na K Rb Cs Fr
@@
Ref. 1 --- - --- - $164.11$ $162.7\pm 0.5$ $289.7\pm 0.3$ $319.8\pm 0.3$ $400.9\pm 0.7$ $317.8\pm 2.4$
Group 2 Be Mg Ca Sr Ba Ra